Marc Horowitz

24.151

February 28, 1992

(1) The sense of ‘the sense of ‘the morning star’ ’ is the same as the sense of ‘the morning star’.

We know that two signs with the same sense have the same reference. If (1) is true, then the reference of ‘the sense of ‘the morning star’ ’ and the reference of ‘the morning star’ are the same. The reference of ‘the morning star’ is the planet Venus. The reference of ‘the sense of ‘the morning star’ ’ is a sense. Senses are objects, but they are not planets. Therefore, (1) is not true. This is not a terribly consequential statement, except that it supports the observation that senses are not physical objects.

I object to Frege’s notion of sense. It seems to me that senses are an artificial construction. I believe in signs, since I can write one down. I believe in references, since they are observable all around me. But under Frege’s theory, senses completely define a sign, making the reference unnecessary. But references exist. This leads me to believe that senses are not fundamental, but could, with effort, be derived somehow from signs and references.